# Latent Neural Operator for Solving Forward and Inverse PDE Problems

### v1

@@ Abstract
Neural operators effectively solve PDE problems from data without knowing the explicit equations, which learn the map from the input sequences of observed samples to the predicted values. Most existed works build the model in the original geometric space, leading to high computational costs when the number of sample points is large. We present the Latent Neural Operator (LNO) solving PDEs in the latent space. In particular, we first propose Physics-Cross-Attention (PhCA) transforming representation from the geometric space to the latent space, then learn the operator in the latent space, and finally recover the real-world geometric space via the inverse PhCA map. Our model retains flexibility that can decode values in any position not limited to locations defined in training set, and therefore can naturally perform interpolation and extrapolation tasks particularly useful for inverse problems. Moreover, the proposed LNO improves in both prediction accuracy and computational efficiency. Experiments show that LNO reduces the GPU memory by 50%, speeds up training 1.8 times, and reaches state-of-the-art accuracy on four out of six benchmarks for forward problems and a benchmark for inverse problem.
@@ 1 Introduction
Neural network approaches for partial differential equations (PDEs) attract increasing attentions in diverse scientific fields, for instance, meteorological prediction[1], industrial design[2], geological sensing[3] and environmental monitoring[4] to name a few. Compared with traditional numerical methods, e.g., finite element method[5], which demands large amount of computational resources and requires meticulous discretization involving a lot of specialized knowledge, neural networks provide a new paradigm for solving PDE-driven problems with better trade-off between accuracy and flexibility.
@@
Recently, Transformers[16] are prevailing in neural operator structures. Theoretically, attention mechanism was proved to be a special form of the integral kernel[15, 17] in neural operator, which naturally conforms to sequence-to-sequence characterization. Moreover, the fully-connected attention structure enables the model characterizes long-distance and quadratic relationship among sample points, which yields more accurate prediction than the vanilla MLP structure[13] but at an expense of increasing the computational complexity drastically, as the complexity is quadratic respect to the length of input sequence.
To alleviate the complexity, existing works employed linear attention mechanism [18, 19, 20, 21] speeding up the computation but sacrificing precision due to their limited expressive power. Another lines of works attempted to solve the PDE in latent space with small number of samples [22, 23, 24], which get rid of abundant sample points in the original geometric space and capture physical correlations in the tighter latent space.
@@
The proposed PhCA module decouples the locations of input samples fed in the encoder and output samples fed in the decoder, which can predict values in any position not limited by the observation. Therefore, it can naturally perform interpolation and extrapolation tasks, which is fundamental for solving inverse problems. The proposed LNO improves in both prediction accuracy and computational efficiency. Experiments shows that LNO reduces the GPU memory by 50% and speeds up training 1.8 times. Moreover, it reaches state-of-the-art accuracy on four (Pipe, Elasticity, Darcy, Navier-Stokes) out of six benchmarks for forward problems and a benchmark for inverse problem on Burgers’ equation.
Finally, the main contributions of our LNO framework are summarized as follow.
  • Flexibility: We propose Physics-Cross-Attention (PhCA) module which decouples the locations of input and output samples and learns the latent space from data. We also build the Latent Neural Operator (LNO) model for solving both forward and inverse PDE problems.
  • Efficiency: The LNO model transforms the data to the latent space only once. Compared with other approaches that switch the latent space and geometric space in each Transformer layer, the proposed model reduces complexity drastically in term of GPU memory usage, training time and number of model parameters.
@@ 2.1 Physics-Informed Machine Learning
Physics-informed machine learning methods incorporate prior knowledge of PDEs into loss function or model architecture. The Physics-Informed Neural Network (PINN)[7] and its variants are among the most classic methods. PINN integrates the differential equation, initial conditions and boundary conditions into the loss function, guiding the neural network to approximate the solution by ensuring its derivatives satisfies the equation and the output aligns with the initial and boundary conditions. Other methods attempt to encode physical priors into the model architecture, e.g., the Finite Element Neural Network (FINN)[8] and Physics-encoded Recurrent Convolutional Neural Network (PeRCNN)[10]. These methods achieve high precision, but require solving optimization to train the networks for each instance. In addition, in the case, where concrete PDE is unknown or not exact, it is challenging to adopt those methods, limiting the generalization ability and flexibility in practical data-rich situations.
@@ 2.2 Operator Learning
A series of kernel integral operator forms were proposed to model the mappings between functions as neural operators[26, 14, 15]. Specifically, Galerkin Transformer[20] utilizes Galerkin-type attention mechanism as kernel integral operators. OFormer[27] improves Galerkin-type attention by decoupling observation positions and query positions through cross attention. GNOT[21] introduces heterogeneous cross attention to modulate various input conditions for query positions, and alleviates computational pressure of the large grid by linear-time-complexity attention mechanism. FactFormer[28] proposes axial factorized kernel integral, decomposing input functions into multiple one-dimensional sub-functions, effectively reducing computation. ONO[29] enhances the generalization of neural operators by adding orthogonality regularization in attention mechanisms.
Researchers also explored the idea of modeling in latent space. Transolver[23] employs Physics-Attention in each layer to map back-and-forth between geometric features and physical features. LSM[22] encodes the input function into the latent space using cross-attention, constructs orthogonal basis, and then decodes the solution back to the geometric space through cross-attention again. UPT[24] introduces encoding and decoding losses, compresses geometric point information into super-node information in latent space via graph neural network[30].
Compared with existing methods, we autonomously learn mappings between functions in real-world space and the latent space in an end-to-end fashion with the cross-attention mechanism, without manually constructed orthogonal basis nor additionally introduced loss.
@@ 2.3 Inverse Problem
Theoretical works on solving inverse problems have also been proposed. The reconstruction of diffusion fields with localized sources have been studied[35] from the perspective of sampling matrix in spatiotemporal domain. PINN[36, 37, 38] approximates the solution based on partially observed data to compute the unknown data. NIO[39] combines DeepONet[13] and FNO[14] to construct a framework for solving inverse problems. FINN[8, 40] sets boundary conditions as learnable parameters, and infer boundary values through backpropagation. We aim to unify the solution of forward and inverse problems in the latent space through the PhCA module, which decouples observation and prediction positions.
@@ 3 Method
We formally define both the forward and the inverse problems of PDEs. Then, we introduce our Latent Neural Operator (LNO) model. Finally, we discuss the core module that learns transformation between real-world space and latent space via cross-attention mechanism.
@@ 3.1 Problem Setup
We consider a partial differential equation (PDE) with boundary and/or initial conditions defined on $D\subseteq\mathbb{R}^{d}$ or $\mathbb{R}^{d}\times[0,T]$
@@
$u({x})$ $=b({x}),\quad{x}\in\partial D$
where $\mathcal{L}_{a}$ is an operator containing partial differentials parameterized by a set of coefficients $a$, and $u(x)$, $x\in D$ and $b({x})$ represent the solution to the PDE and the boundary/initial conditions respectively.
A conventional forward problem aims to find the solution $u(x)$, given the differential operator $\mathcal{L}_{a}$ in (1) with the explicit parameters $a$ and the boundary/initial conditions $b(x)$. For an inverse problem, given an observed set of partial data $\{(x,u(x))|\;x\in\tilde{D}\subset D\}$, the task is to infer the PDE that generates the observed data along with possible boundary/initial conditions or the parameters $a$ in the differential operator $\mathcal{L}_{a}$.
Both forward and inverse problems can be unified as an operator learning task fitting into a sequence-to-sequence translation framework. An operator model estimates the mapping $\mathcal{F}:h\mapsto g$ from an input function $h$ to an output function $g$ by the data-driven approach using paired training data of input-output sequences ($(\{h(y_{i})\}_{i},\{g(z_{j})\}_{j})$.
Specifically, for the forward problem, the inputs are samples of either boundary/initial function $b$ or PDE coefficient function $a$, and the output are the physical quantity $u$ evaluated at a grid of points. Conversely, for the inverse problem, the input are a few partially observation samples $\{(pos_{in}^{(i)},val_{in}^{(i)})\}^{N_{in}}=\{(x_{i},u(x_{i}))\}_{i=1}^{N_{in}}$, and the outputs are boundary/initial function (e.g., infer initial state) or the PDE coefficient function (also coined as system identification problem).
The challenge in solving forward and inverse problems of PDEs lies in the large lengths of input and output sequences. In terms of operator approximating, complex kernel integrals lead to high computational cost while simple kernel integrals lack in accuracy. In this work, we build accurate and computationally efficient model for both forward and inverse problem by learning the operator in the learnable latent space.
@@ 3.2 Latent Neural Operator
The proposed model of Latent Neural Operator (LNO) is applicable for both forward and inverse PDE problems. It consists of four modules: an embedding layer to lift the dimension of the input data, an encoder to transform the input data to a learnable latent space, a sequence of self-attention layers for modeling operator in the latent space, and a decoder to recover the latent representation to the real-world space. The overall architecture of LNO is shown in Figure 1
@@
First, we embed the positions and physics quantities of the input sequence to a higher dimensional representation. It is implemented by two MLPs, where the first only embeds the position information and the second embeds both the location and physics quantities
${x}^{(i)}$ $={\rm trunk\text{-}projector}\big{(}{pos}_{in}^{(i)}\big{)},$ ${x}^{(i)}\in\mathbb{R}^{d\times 1}$
${y}^{(i)}$ $={\rm branch\text{-}projector}\big{(}{\rm concat}({pos}_{in}^{(i)},{val}_{in}^{(i)})\big{)},\quad$ ${y}^{(i)}\in\mathbb{R}^{d\times 1}.$
Analogous to DeepONet [13], we named these MLPs as trunk-projector and branch-projector respectively. The embedding operation helps to map the locations and physics quantities to the space where their relationships are easier to be captured.
Encoding to latent space
We seek the representation of the input function in a latent space, where the number $M$ of sample points in the latent space can be much smaller than the number $N_{in}$ of samples of raw inputs. We use the cross-attention to model space transformation, where the query $\boldsymbol{H}$, key $\boldsymbol{X}=[x^{(1)},x^{(2)},\ldots,x^{(N_{in})}]^{T}$ and value $\boldsymbol{Y}=[y^{(1)},y^{(2)},\ldots,y^{(N_{in})}]^{T}$ inputs corresponding to locations in the target latent space, embedding of locations in the source real-world space, and embedding of locations and physics quantities in the source real-world respectively
$\boldsymbol{Z}^{0}={\rm Physics\text{-}Cross\text{-}Attention}(\boldsymbol{H},\boldsymbol{X},\boldsymbol{Y})$ (2)
In contrast to FNO[14], where the latent space is pre-defined as the frequency domain, we learn the latent space from data. Therefore, the sample locations $\boldsymbol{H}$ as query in the cross-attention are also learnable parameters.
After encoding the input function into the latent space, the length of the sequence to be processed is reduced from $N_{in}$ to $M$. With $M\ll N_{in}$, extracting the feature of the input function in the latent space will be much more efficient than in the real-world geometry space.
@@
We model the operator of the forward/inverse PDE problem in the latent space, using a stack of Transformer blocks with self-attention mechanism
@@
where $l\in\{1,2,...,L\}$ is the index of Transformer block with $L$ being the number of the blocks. We experiment several implementations of attention mechanisms in the ablation study, and find the classical scaled dot-product attention [16] performs consistently well on various tasks.
Decoding to real-world space
Finally, we decode the output sequence from the latent sequence through Physics-Cross-Attention according to the query ${pos}_{out}$ of the output sample locations, and use another MLP to map the embbedings to the values of output function
@@
During the inferring phase, the trunk-projector can accept inputs from any position, enabling LNO generalize to unseen region where the positions are not presented in training phase.
3.3 Physics Cross Attention
The Physics-Cross-Attension (PhCA) is the core module in the aforementioned latent neural operator (LNO) model, acted as the encoder and decoder which transform the data representation back-and-forth between the real-world space and the latent space. Motivated by the work [21] that learns the operator mapping cross domains using the cross-attention, we use the same cross-attention to model the transformation to/from the latent space, but set the latent space learnable rather than pre-defined, e.g., FNO [14].
In particular, the rows of the query $\boldsymbol{H}$ and key $\boldsymbol{X}$ matrices in (2) represent the embedded positions of samples in the target and source space respectively. A row of the value matrix $\boldsymbol{Y}$ is the embedding of both position and physics quantity of a sample in the source space. We design the latent space to be learned from data. Thus, the positions $\boldsymbol{H}$ of samples in the latent space should be set as learnable parameters as well. Then, the cross-attention of the encoder in (2) is simplified as
$\boldsymbol{Z}^{0}={\rm softmax}\Big{(}\tfrac{1}{\sqrt{d}}\boldsymbol{H}\boldsymbol{W}_{q}\boldsymbol{W}_{k}^{T}\boldsymbol{X}^{T}\Big{)}\boldsymbol{Y}\boldsymbol{W}_{v}={\rm softmax}\big{(}\boldsymbol{W}_{1}\boldsymbol{X}^{T}\big{)}\boldsymbol{Y}\boldsymbol{W}_{v},$ (4)

### v3

@@ Abstract
Neural operators effectively solve PDE problems from data without knowing the explicit equations, which learn the map from the input sequences of observed samples to the predicted values. Most existing works build the model in the original geometric space, leading to high computational costs when the number of sample points is large. We present the Latent Neural Operator (LNO) solving PDEs in the latent space. In particular, we first propose Physics-Cross-Attention (PhCA) transforming representation from the geometric space to the latent space, then learn the operator in the latent space, and finally recover the real-world geometric space via the inverse PhCA map. Our model retains flexibility that can decode values in any position not limited to locations defined in the training set, and therefore can naturally perform interpolation and extrapolation tasks particularly useful for inverse problems. Moreover, the proposed LNO improves both prediction accuracy and computational efficiency. Experiments show that LNO reduces the GPU memory by 50%, speeds up training 1.8 times, and reaches state-of-the-art accuracy on four out of six benchmarks for forward problems and a benchmark for inverse problem. Code is available at https://github.com/L-I-M-I-T/LatentNeuralOperator.
@@ 1 Introduction
Neural network approaches for partial differential equations (PDEs) attract increasing attention in diverse scientific fields, for instance, meteorological prediction[1], industrial design[2], geological sensing[3] and environmental monitoring[4] to name a few. Compared with traditional numerical methods, e.g., finite element method[5], which demands a large number of computational resources and requires meticulous discretization involving a lot of specialized knowledge, neural networks provide a new paradigm for solving PDE-driven problems with a better trade-off between accuracy and flexibility.
@@
Recently, Transformers[16] prevail in neural operator structures. Theoretically, the attention mechanism was proved to be a special form of the integral kernel[15, 17] in neural operator, which naturally conforms to sequence-to-sequence characterization. Moreover, the fully-connected attention structure enables the model to characterize long-distance and quadratic relationships among sample points, which yields more accurate prediction than the vanilla MLP structure[13] but at the expense of increasing the computational complexity drastically, as the complexity is quadratic respect to the length of input sequence.
To alleviate the complexity, existing works employed linear attention mechanism [18, 19, 20, 21] speeding up the computation but sacrificing precision due to their limited expressive power. Another line of works attempted to solve the PDE in the latent space with a small number of samples [22, 23, 24, 25, 26], which get rid of abundant sample points in the original geometric space and capture physical correlations in the tighter latent space.
@@
The proposed PhCA module decouples the locations of input samples fed in the encoder and output samples fed in the decoder, which can predict values in any position not limited by the observation. Therefore, it can naturally perform interpolation and extrapolation tasks, which is fundamental for solving inverse problems. The proposed LNO improves both prediction accuracy and computational efficiency. Experiments show that LNO reduces the GPU memory by 50% and speeds up training 1.8 times. Moreover, it reaches state-of-the-art accuracy on four (Pipe, Elasticity, Darcy, Navier-Stokes) out of six benchmarks for forward problems and a benchmark for inverse problem on Burgers’ equation.
The main contributions of our LNO framework are summarized as follows.
  • Flexibility: We propose the Physics-Cross-Attention (PhCA) module which decouples the locations of input and output samples and learns the latent space from data. We also build the Latent Neural Operator (LNO) model for solving both forward and inverse PDE problems.
  • Efficiency: The LNO model transforms the data to the latent space only once. Compared with other approaches that switch the latent space and geometric space in each Transformer layer, the proposed model reduces complexity drastically in terms of GPU memory usage, training time and number of model parameters.
@@ 2.1 Physics-Informed Machine Learning
Physics-informed machine learning methods incorporate prior knowledge of PDEs into loss function or model architecture. The Physics-Informed Neural Network (PINN)[7] and its variants are among the most classic methods. PINN integrates the differential equation, initial conditions and boundary conditions into the loss function, guiding the neural network to approximate the solution by ensuring its derivatives satisfies the equation and the output aligns with the initial and boundary conditions. Other methods attempt to encode physical priors into the model architecture, e.g., the FInite volume Neural Network (FINN)[8] and Physics-encoded Recurrent Convolutional Neural Network (PeRCNN)[10]. These methods achieve high precision but require solving optimization to train the networks for each instance. In addition, in the case, where concrete PDE is unknown or not exact, it is challenging to adopt those methods, limiting the generalization ability and flexibility in practical data-rich situations.
@@ 2.2 Operator Learning
A series of kernel integral operator forms were proposed to model the mappings between functions as neural operators[28, 14, 15]. Specifically, Galerkin Transformer[20] utilizes the Galerkin-type attention mechanism as kernel integral operators. OFormer[29] improves Galerkin-type attention by decoupling observation positions and query positions through cross-attention. GNOT[21] introduces heterogeneous normalized cross-attention to modulate various input conditions for query positions and alleviates computational pressure of the large grid by linear-time-complexity attention mechanism. FactFormer[30] proposes axial factorized kernel integral, decomposing input functions into multiple one-dimensional sub-functions, effectively reducing computation. ONO[31] enhances the generalization of neural operators by adding orthogonality regularization in attention mechanisms.
Researchers also explored the idea of modeling in the latent space. Transolver[23] employs Physics-Attention in each layer to map back-and-forth between geometric features and physical features. LSM[22] encodes the input function into the latent space using cross-attention, constructs a set of orthogonal bases, and then decodes the solution back to the geometric space through cross-attention again. UPT[24] introduces encoding and decoding losses, and compresses geometric point information into super-node information in the latent space via graph neural network[32].
Compared with existing methods, we autonomously learn mappings between functions in the real-world space and the latent space in an end-to-end fashion with the cross-attention mechanism, without manually constructed orthogonal basis nor additionally introduced loss.
@@ 2.3 Inverse Problem
Theoretical works on solving inverse problems have also been proposed. The reconstruction of diffusion fields with localized sources has been studied[37] from the perspective of sampling matrix in the spatiotemporal domain. PINN[38, 39, 40] approximates the solution based on partially observed data to compute the unknown data. NIO[41] combines DeepONet[13] and FNO[14] to construct a framework for solving inverse problems. FINN[8, 42] sets boundary conditions as learnable parameters, and infers boundary values through backpropagation. The generative method[43] generates the initial conditions through latent codes and refines them based on the results of time evolution. We aim to unify the solution of forward and inverse problems in the latent space through the PhCA module, which decouples observation and prediction positions.
@@ 3 Method
We formally define both the forward and the inverse problems of PDEs. Then, we introduce our Latent Neural Operator (LNO) model. Finally, we discuss the core module that learns the transformation between the real-world geometric space and the latent space via the cross-attention mechanism.
@@ 3.1 Problem Setup
We consider a partial differential equation (PDE) with boundary and/or initial conditions defined on $\Omega\subseteq\mathbb{R}^{d}$ or $\mathbb{R}^{d}\times[0,T)$
@@
$u({x})$ $=b({x}),\quad{x}\in\partial\Omega$
where $\mathcal{L}_{a}$ is an operator containing partial differentials parameterized by a set of coefficients $a$, and $u(x)$, $x\in\Omega$ and $b({x})$ represent the solution to the PDE and the boundary/initial conditions respectively.
A conventional forward problem aims to find the solution $u(x)$, given the differential operator $\mathcal{L}_{a}$ in (1) with the explicit parameters $a$ and the boundary/initial conditions $b(x)$. For an inverse problem, given an observed set of partial data $\{(x,u(x))|\;x\in\tilde{\Omega}\subset\Omega\}$, the task is to infer the PDE that generates the observed data along with possible boundary/initial conditions or the parameters $a$ in the differential operator $\mathcal{L}_{a}$.
Both forward and inverse problems can be unified as an operator learning task fitting into a sequence-to-sequence translation framework. An operator model estimates the mapping $\mathcal{F}:f\mapsto g$ from an input function $f$ to an output function $g$ by the data-driven approach using paired training data of input-output sequences ($(\{f(y_{i})\}_{i},\{g(z_{j})\}_{j})$, where $y_{i}$ and $z_{j}$ correspond to positions in the domain of the input and output function respectively.
Specifically, for the forward problem, the inputs are samples of either boundary/initial function $b$ or PDE coefficient function $a$, and the output is the physical quantity $u$ evaluated at a grid of points. Conversely, for the inverse problem, the input is a few partially observation samples $\{(pos_{in}^{(i)},val_{in}^{(i)})\}^{N_{in}}_{i=1}=\{(x_{i},u(x_{i}))\}_{i=1}^{N_{in}}$, and the outputs are boundary/initial function (e.g., infer initial state) or the PDE coefficient function (also coined as system identification problem).
The challenge in solving forward and inverse problems of PDEs lies in the large lengths of input and output sequences. In terms of operator approximating, complex kernel integrals lead to high computational cost while simple kernel integrals lack accuracy. In this work, we build an accurate and computationally efficient model for both forward and inverse problems by learning the operator in the learnable latent space.
@@ 3.2 Latent Neural Operator
The proposed model of Latent Neural Operator (LNO) is applicable for both forward and inverse PDE problems. It consists of four modules: an embedding layer to lift the dimension of the input data, an encoder to transform the input data to a learnable latent space, a series of self-attention layers for modeling operator in the latent space, and a decoder to recover the latent representation to the real-world geometric space. The overall architecture of LNO is shown in Figure 1.
@@
The geometric space is the original space of the PDE input or output which contains $N_{in}$ or $N_{out}$ samples, each with $d$-dimensional position coordinates (which may additionally include a $1$-dimensional time coordinate for time-dependent PDEs) and $n$-dimensional physical quantity values.
First, we embed the positions and physics quantity values of the input sequence in the geometric space to higher $D$-dimensional representations. This process is implemented by two MLPs, where the first only embeds the position information and the second embeds both the location and physics quantity information
${\hat{x}}^{(i)}$ $={\rm trunk\text{-}projector}\big{(}{pos}_{in}^{(i)}\big{)},$ ${\hat{x}}^{(i)}\in\mathbb{R}^{D\times 1}$
${\hat{y}}^{(i)}$ $={\rm branch\text{-}projector}\big{(}{\rm concat}({pos}_{in}^{(i)},{val}_{in}^{(i)})\big{)},\quad$ ${\hat{y}}^{(i)}\in\mathbb{R}^{D\times 1}.$
Analogous to DeepONet [13], we name these MLPs as trunk-projector and branch-projector respectively. The embedding operation helps to map the locations and physics quantities to the embedding space where their relationships are easier to capture.
Removing the physics quantity values from the trunk-projector input grants our model the decoupling property. In this way, we can predict values at positions not included in the sampled input function from the latent space, thus achieving the decoupling of observation and prediction positions. Since the sampled input function always includes position-value pairs, using both position and value information in the branch projector will not affect the decoupling objective.
Encoding to the latent space
We attempt to represent the input function using the representation tokens of $M$ hypothetical sampling positions $\{h^{(k)}\}^{M}_{k=1},h^{(k)}\in R^{D\times 1}$ which exist in the latent space, where $M$ can be much smaller than the number $N_{in}$ of samples of raw inputs.
We use the Physics-Cross-Attention to model the transformation from the geometric space to the latent space. The inputs are query $\boldsymbol{H}=[h^{(1)},h^{(2)}...,h^{(M)}]^{T}$, key $\hat{\boldsymbol{X}}=[\hat{x}^{(1)},\hat{x}^{(2)},\ldots,\hat{x}^{(N_{in})}]^{T}$ and value $\hat{\boldsymbol{Y}}=[\hat{y}^{(1)},\hat{y}^{(2)},\ldots,\hat{y}^{(N_{in})}]^{T}$ which corresponding to embedding of hypothetical locations in the target latent space, embedding of locations in the source real-world geometric space, and embedding of the concatenation of locations and physics quantities in the source real-world geometric space respectively. The output can be calculated through
$\boldsymbol{Z}^{0}={\rm Physics\text{-}Cross\text{-}Attention}(\boldsymbol{H},\boldsymbol{\hat{X}},\boldsymbol{\hat{Y}})$ (2)
In contrast to FNO[14], where the latent space is pre-defined as the frequency domain, we learn the latent space from data. Therefore, the embeddings of sample locations $\boldsymbol{H}$ as queries in the cross-attention are also learnable parameters.
After encoding the input function into the latent space, the length of the sequence to be processed is reduced from $N_{in}$ to $M$. With $M\ll N_{in}$, extracting and converting the feature of the input function in the latent space will be much more efficient than in the real-world geometric space.
@@
We model the operator of the forward/inverse PDE problem in the latent space to convert the feature of the input function to that of the output function, using a stack of Transformer blocks with the self-attention mechanism
@@
where $l\in\{1,2,...,L\}$ is the index of Transformer block with $L$ being the total number of the blocks. We experiment with several implementations of attention mechanisms in the ablation study and find the classical scaled dot-product attention[16] performs consistently well on various tasks.
Decoding to the geometric space
Finally, we decode the output sequence from the latent sequence through Physics-Cross-Attention according to the query ${pos}_{out}$ of the output sampling locations and use another MLP to map the embeddings to the values of the output function
@@
where $\boldsymbol{P}=[p^{(1)},p^{(2)},\ldots,p^{(N_{out})}]^{T}$ and $\boldsymbol{U}=[u^{(1)},u^{(2)},\ldots,u^{(N_{out})}]^{T}$. During the inferring phase, the trunk-projector can accept inputs from any position, enabling LNO to generalize to the unseen region where the positions are not presented in the training phase.
3.3 Physics-Cross-Attention
The Physics-Cross-Attension (PhCA) is the core module in the aforementioned latent neural operator (LNO) model, acting as the encoder and decoder which transform the data representation back-and-forth between the real-world geometric space and the latent space. Motivated by the work [22] that learns the operator mapping across domains using the cross-attention, we use the same cross-attention to model the transformation to/from the latent space but set the latent space learnable rather than pre-defined, e.g., triangular space or frequency domain [14].
In particular, the rows of the query $\boldsymbol{H}$ and key $\hat{\boldsymbol{X}}$ matrices in (2) represent the embedded positions of samples in the target and source space respectively. A row of the value matrix $\hat{\boldsymbol{Y}}$ is the embedding of both the position and physics quantity of a sample in the source space. We design the latent space to be learned from data. Thus, the positions $\boldsymbol{H}$ of samples in the latent space should be set as learnable parameters as well. Then, the cross-attention of the encoder in (2) is simplified as
$\boldsymbol{Z}^{0}={\rm softmax}\Big{(}\tfrac{1}{\sqrt{D}}\boldsymbol{H}\boldsymbol{W}_{q}\boldsymbol{W}_{k}^{T}\boldsymbol{\hat{X}}^{T}\Big{)}\boldsymbol{\hat{Y}}\boldsymbol{W}_{v}={\rm softmax}\big{(}\boldsymbol{W}_{1}\boldsymbol{\hat{X}}^{T}\big{)}\boldsymbol{\hat{Y}}\boldsymbol{W}_{v},$ (4)